OpenStudy (anonymous):

Complete the square, solve the equation for x. x^2=24-4x. The answer is x= -2 ± 2√(7)

OpenStudy (anonymous):

Step I: Rewrite your equation as x^2 + 4x + 24

OpenStudy (anonymous):

Now take 1/2 of the b term (the 4) and square it. Add that number to both sides of the equation. So 1/2 of 4 is 2; and 2 squared is 4. We will add 4 to both sides of the equation, you get x^2 + 4x + 4 = 24 + 4

OpenStudy (anonymous):

The left side is a perfect square, so you have (x + 2)^2 = 28

OpenStudy (anonymous):

If (x + 2)^2 = 28, then x+2 = +/- sqrt (28)

OpenStudy (anonymous):

so x = - 2 +/- sqrt (28)

OpenStudy (anonymous):

Sqrt (28) = sqrt (4) times sqrt (7) = 2 sqrt (7)

OpenStudy (anonymous):

So your final answer is - 2 +/- 2 sqrt (7)

OpenStudy (anonymous):

Done.

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

Welcome.

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

Incorrect. Watch, I will show you.

OpenStudy (anonymous):

2x^2 + 8x - 15 = 0 The very first step is to add 15 to both sides! You have 2x^2 + 8x = 15

OpenStudy (anonymous):

Now, you CANNOT take 1/2 of the b term and square it UNLESS you have the coefficient of x^2 being 1. We have 2x^2. So FIRST, divide both sides by 2.

OpenStudy (anonymous):

So we have x^2 + 4x = ,15/2

OpenStudy (anonymous):

Now, that the coefficient of x^2 is 1, we NOW take 1/2 of the b term and square it. And we will add that number to both sides. So 1/2 of 4 is 2, and 2 squared is 4. We will add 4 to both sides.

OpenStudy (anonymous):

So we have x^2 + 4x + 4 = 15/2 + 4 or x^2 + 4x + 4 = 11.5

OpenStudy (anonymous):

wait where did 4x come from? it was 8x.

OpenStudy (anonymous):

so x^2+8x+16=15/2+16 ?

OpenStudy (anonymous):

We had 2x^2 + 8x on the left side, all what I did was divide both sides by 2, so you get x^2 + 4x.

OpenStudy (anonymous):

OH i forgot about that step. okay that makes sense

OpenStudy (anonymous):

So what I am saying is that if the left side was 3x^2 + 13x, we would have to first divide the left side by 3 and the right side by 3. What is critical is the coefficient of x^2 be 1. Thats a must.

OpenStudy (anonymous):

So we have x^2 + 4x + 4 = 11.5

OpenStudy (anonymous):

The left side is a perfect square, it is (x+2)^2

OpenStudy (anonymous):

So, we have (x+2)^2 = 11.5, so x+ 2 = +/- sqrt (11.5) so x = -2 +/- sqrt (11.5)

OpenStudy (anonymous):

Done.

OpenStudy (anonymous):

Make sure that each sense is crystal clear. Any questions, just write to me.

OpenStudy (anonymous):

all of that makes sense to me. but the book has the answers in the back, thats why i have to show work, but it says the answer is x=-2±\[\left(\begin{matrix}\sqrt{46} \\ 2\end{matrix}\right)\]